Ticket Number: HEA-NOC/20080225-10 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080225 14:32 UTC Problem Start: 20080228 08:00 UTC
Ticket Update: 20080228 08:20 UTC Problem End: 20080228 09:00 UTC
Site/line: IT Carlow
Migration of IT Carlow's current connections to HEAnet to new IP
equipment in Citywest and Kilcarbery.
No service interruption is expected, BGP sessions need to be reset and
will be done gradually. A side effect of this migration is asymmetrical
routing while convergence takes place.
Once the changes are completed, traffic will use the new connections,
continuing to have the current ones as backup. Eventually these will be
Contingency plan is to revert changes to existing setup by disabling any
routing over the new links.
20080225 14:30 GL Opening ticket. Confirmed with IT Carlow that the
will take place on Thursday 28th at 8am. Configurations in place.
20080228 08:20 GL Migration completed. Traffic now routing through CRS-1
Citywest. Failover testing performed. Successful to CRS-1 in Kilcarbery.
connections still in place and will be decommissioned in the coming weeks.
monitor for 24 hours
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080225-10
HEAnet Limited. Registered in Ireland, No. 275301.